What is the volume of a sample of ethane at 467k and 1.1atm ,if it occupies 405ml at 298k and 1.1atm


1
Expert's answer
2022-03-03T13:26:03-0500

According to the combined gas law,


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


Therefore,


"V_2=\frac{P_1V_1T_2}{T_1P_2}=\frac{1.1\ atm\times405\ mL\times467\ K}{298\ K\times1.1\ atm}=635\ mL"


Answer: 635 mL